OpenStudy (anonymous):

Use the quadratic formula to find the zeros of the function. Round to tenths if necessary. y = 3x2 + 5x A. {-2.23, 1.7} B. {-1.7, 0} C. {-2.8, 1.7} D. no real solution

hero (hero):

a = 3 b = 5 c = 0 \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] \[x = \frac{-5 \pm \sqrt{5^2 - 4(3)(0)}}{2(3)}\] \[x = \frac{-5 \pm \sqrt{5^2}}{2(3)}\]

hero (hero):

You may be able to finish solving from there.

OpenStudy (anonymous):

so no real solution?

hero (hero):

Did you forget the c value when you were writing the quadratic equation?

hero (hero):

The solution according to what you posted is (0,0)

OpenStudy (anonymous):

so 0,-1.7

hero (hero):

Which is not in the answer choices...

hero (hero):

No, that is not a solution

hero (hero):

y = 3(0)^2 + 5(0) = 0 + 0 = 0

hero (hero):

(0,0) works (0,-1.7) does not work

hero (hero):

(-1.7,0) doesn't work either

hero (hero):

In fact, none of the given answer choices work

hero (hero):

I'm pretty certain you wrote the problem incorrectly to begin with because the answer choices do not match the question.

OpenStudy (anonymous):

\[-5\pm \sqrt{25-0}/6 \] thats the equation i get

hero (hero):

Once again, the answer choices do not match the problem. Double check to make sure you have posted it from your book correctly.

OpenStudy (anonymous):

y = 3x2 + 5x

OpenStudy (anonymous):

the 2 is a exponent

hero (hero):

You're telling me something I already know.

hero (hero):

You insert the coefficients a, b, and c into the quadratic formula.

hero (hero):

But you don't seem to have a c value. A quadratic equation in general form is \(ax^2 + bx + c = 0\)

hero (hero):

You have \(3x^2 + 5x = 0\) but you don't have a \(c\) value.

OpenStudy (anonymous):

the question asks to fine the 0

OpenStudy (anonymous):

find

hero (hero):

I found the zero.

OpenStudy (anonymous):

Use the quadratic formula to find the zeros of the function. Round to tenths if necessary

hero (hero):

x = 0 and y = 0 I already demonstrated it to you above.

hero (hero):

There is something about what I did that you don't understand apparently.

OpenStudy (anonymous):

i get it. but your not giving a answer. so i dont get what your talking about now